OpenStudy (anonymous):

(a)What is the wavelength (in nanometers) of light having the frequency of 8.6x10^13 Hz? (b) What is the frequency (in Hz) of light having a wavelength of 566nm?

OpenStudy (anonymous):

a) lambda (wavelenght) = c/f = 3*10^8/8,6*10^13 = 348 nm b) 5,3*10^13 Hz

sam (.sam.):

\[v=f \lambda\]